OpenStudy (anonymous):

^3√54 + ^3√16 - ^3√24 Answers I can choose from: a. ^3√46 b. 4 c. 5^3√2 - 2^3√3 d. 4^3√2

OpenStudy (anonymous):

I'm assuming those are cube roots. All you need to do is add up everything inside the roots. A good way to understand it is if you set the equation equal to x: \[\sqrt[3]{54}+\sqrt[3]{16}-\sqrt[3]{24} = x\] Then if you raise all the terms to the power of 3 you end up with: \[54+16-24=x^3 \] Once you have that you just add up the numbers on the left and you get \[46=x^3\] Then you take the cube root of both sides again to figure out x which ends up being: \[\sqrt[3]{46}=x\] Which is answer A

OpenStudy (anonymous):

Thank you so much! (: